Skip to main contentSkip to solution

There is a 4-in-5 chance that A can complete the marathon. There is a 5-in-8 chance that B can't complete the marathon, which is independent of A's performance. What is the probability that both A and B end up completing?

Solution

✅ Correct Option: 3

Probability A completes marathon =45= \dfrac{4}{5}

Probability B can't complete marathon =58= \dfrac{5}{8}

A's and B's performances are independent.


The problem gives the probability that B can't complete.

P(B can’t complete)=58P(B \text{ can't complete}) = \dfrac{5}{8}

P(B can complete)=1−58P(B \text{ can complete}) = 1 - \dfrac{5}{8}

P(B can complete)=38P(B \text{ can complete}) = \dfrac{3}{8}


Since A and B perform independently:

P(Both complete)=P(A completes)×P(B completes)P(\text{Both complete}) = P(A \text{ completes}) \times P(B \text{ completes})

P(Both complete)=45×38P(\text{Both complete}) = \dfrac{4}{5} \times \dfrac{3}{8}

P(Both complete)=1240P(\text{Both complete}) = \dfrac{12}{40}

P(Both complete)=310P(\text{Both complete}) = \dfrac{3}{10}

Therefore, the probability that both A and B complete the marathon is 310\dfrac{3}{10}.

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question